A raffle has a first prize of $400, two second prizes of $ 75 eac, and ten third prizes of $20 each. One thousand tickets are sold at $1 a piece. Find the expected winnings of a person buying a ticket.
Answer by solver91311(24713)



First, how much would it cost to buy all of the tickets? 1000 X 1 = $1000.

How much would you win if you won all of the prizes? 400 + 2*75 + 10*20 = 400 + 300 + 200 = 900.

900 divided by 1000 is 0.9 and 0.9 times $1 is 90 cents. Your expected winnings per ticket are 90 cents. In other words it costs you 1 dollar to win 90 cents. Don't play unless the proceed go for what you consider to be a good cause and you don't mind donating some of your money to that cause.

John